OpenStudy (anonymous):

add or subtract. simplify. (5x)/ (x^2-x-6) - (4) / (x^2+4x+4)

OpenStudy (asnaseer):

the first thing you need to do is to factorise the denominators.

OpenStudy (anonymous):

yes. i did that. i got 5x/ (x-3) (x+2) + 4/ (x+2) (x+2)

OpenStudy (asnaseer):

good - so now you will notice that there is a common factor in both terms in the denominators.

OpenStudy (anonymous):

yeah x+2

OpenStudy (asnaseer):

thats right, so can now simplify this expression by pulling out that common factor as follows:\[\frac{5x}{(x-3)(x+2)}-\frac{4}{(x+2)(x+2)}=\frac{1}{x+2}(\frac{5x}{x-3}-\frac{4}{x+2})\]

OpenStudy (asnaseer):

can you now do the subtraction within the braces?

OpenStudy (anonymous):

so why do we multiply it by 1/x+2?????

OpenStudy (asnaseer):

we pulled out the 1/(x+2) as a common factor of both terms.

OpenStudy (anonymous):

okay.

OpenStudy (asnaseer):

e.g.:\[\frac{5x}{(x-3)(x+2)}=\frac{5x}{x-3}*\frac{1}{x+2}\]

OpenStudy (anonymous):

okay. can you help me multiply within the brackets. i'm still not sure.

OpenStudy (asnaseer):

ok - within the braces, you first need to find the LCM (Lowest Common Multiple) of both denominators. in this case that would be (x-3)(x+2)

OpenStudy (anonymous):

okay. but you know how in the denominator there are 3 of (x+2). do u still just have x+2 be ur least common multiple or do you square it. since it is used 3 times when it is factored

OpenStudy (asnaseer):

then perform the subtraction as follows:\[\frac{5x}{x-3}-\frac{4}{x+2}=\frac{5x}{x-3}*\frac{x+2}{x+2}-\frac{4}{x+2}*\frac{x-3}{x-3}=\frac{5x(x+2)-4(x-3)}{(x+2)(x-3)}\]

OpenStudy (asnaseer):

why do you think there are "3" of (x+2)?

OpenStudy (anonymous):

because when we ffactored both sides of the denominators we got (x-3) (x+2) on one side and (x+2) (x+2) on the other side

OpenStudy (asnaseer):

but we then pulled out the common (x+2) leaving a simpler subtraction within the braces

OpenStudy (anonymous):

okay so when u multiplied it by 1/ x+2 it got rid of one of the x+2. leaving us with only 2 instead of three

OpenStudy (asnaseer):

there is only ONE (x+2) term left within the braces:\[\frac{5x}{(x-3)(x+2)}-\frac{4}{(x+2)(x+2)}=\frac{1}{x+2}(\frac{5x}{x-3}-\frac{4}{x+2})\] it is bit like this:\[\frac{5}{4*3}-\frac{4}{7*3}=\frac{1}{3}(\frac{5}{4}-\frac{4}{7})\]

OpenStudy (asnaseer):

or, even simpler without fractions:\[7*3 - 4*3=3(7-4)\]

OpenStudy (asnaseer):

the "common factor" 3 is pulled out of both terms leaving no 3's within the braces

OpenStudy (anonymous):

okay.

OpenStudy (asnaseer):

right, so I worked out the subtraction within the braces above as:\[\frac{5x}{x-3}-\frac{4}{x+2}=\frac{5x}{x-3}*\frac{x+2}{x+2}-\frac{4}{x+2}*\frac{x-3}{x-3}=\frac{5x(x+2)-4(x-3)}{(x+2)(x-3)}\]

OpenStudy (asnaseer):

do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok - so can you simplify this resulting fraction?

OpenStudy (anonymous):

do you cancel like terms. does it leave us with 5x-4?

OpenStudy (asnaseer):

no, you need to work this out:\[5x(x+2)-4(x-2)=\]

OpenStudy (asnaseer):

sorry that should be 4(x-3)

OpenStudy (asnaseer):

not 4(x-2)

OpenStudy (anonymous):

as 5x^2 +10x-4x+8

OpenStudy (anonymous):

then 5x^2+6x+8

OpenStudy (asnaseer):

^^ 4(x-3) not 4(x-2)

OpenStudy (asnaseer):

\[5x(x+2)-4(x-3)=\]

OpenStudy (anonymous):

okay i got. it . like the way i jsut did it but instead with x-3 not x-2?

OpenStudy (asnaseer):

exactly

OpenStudy (anonymous):

okay thank you for your help.

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

???

OpenStudy (asnaseer):

don't forget to multiply by the common factor we took out at the beginning

OpenStudy (asnaseer):

yw == you're welcome :)

OpenStudy (anonymous):

okay thanks.